# Uncertainty-Aware Hybrid Inference with On-Device Small and Remote Large Language Models

### v1

@@ Abstract
This paper studies a hybrid language model (HLM) architecture that integrates a small language model (SLM) operating on a mobile device with a large language model (LLM) hosted at the base station (BS) of a wireless network. The HLM token generation process follows the speculative inference principle: the SLM’s vocabulary distribution is uploaded to the LLM, which either accepts or rejects it, with rejected tokens being resampled by the LLM. While this approach ensures alignment between the vocabulary distributions of the SLM and LLM, it suffers from low token throughput due to uplink transmission and the computation costs of running both language models. To address this, we propose a novel HLM structure coined Uncertainty-aware HLM (U-HLM), wherein the SLM locally measures its output uncertainty, and skips both uplink transmissions and LLM operations for tokens that are likely to be accepted. This opportunistic skipping is enabled by our empirical finding of a linear correlation between the SLM’s uncertainty and the LLM’s rejection probability. We analytically derive the uncertainty threshold and evaluate its expected risk of rejection. Simulations show that U-HLM reduces uplink transmissions and LLM computation by 45.93%, while achieving up to 97.54% of the LLM’s inference accuracy and 2.54$\times$ faster token throughput than HLM without skipping.
@@ I Introduction
To this end, this paper seeks to find a solution for the HLM framework that improves token throughput while minimizing the degradation of inference accuracy. The approach builds on a straightforward yet effective strategy: skip uplink transmissions as well as LLM computations for tokens that are likely to be accepted. However, implementing this is challenging, as the SLM must independently predict whether a token will be accepted or rejected without access to the LLM’s vocabulary distribution. To enable this capability, we explore a method in which the SLM leverages uncertainty, a measure of the model’s self-assessed confidence in its outputs, to predict the LLM’s rejection probability. Among the various approaches we evaluate for measuring uncertainty in LLMs, our experiments reveal that uncertainty measured through temperature perturbation exhibits a strong linear correlation with the rejection probability in the HLM framework. Building on this observation, we propose a novel HLM framework called Uncertainty-aware HLM (U-HLM). As illustrated in Figure 1, U-HLM enables the device to selectively transmit the SLM’s vocabulary distribution only for tokens with uncertainty exceeding a predefined threshold.
@@ III Uncertainty-Aware Hybrid Language Model for High-Throughput LLM Inference
In this section, we conduct a feasibility experiment inspired by the concept of uncertainty to investigate whether a device can predict the rejection probability of a token on its own without the server’s target probability. Based on this, we propose an Uncertainty-aware Hybrid Language Model (U-HLM), in which the SLM measures the uncertainty and opportunistically skips uplink transmissions to the LLM if the estimated uncertainty does not exceed its threshold. Additionally, we design the uncertainty threshold for U-HLM based on the rejection probability threshold that determines when the device intends to skip transmissions. We also derive the expected rejection risk and its upper bound by incorporating the density of uncertainty into the analysis.
@@ IV-B Inference Accuracy and Token Throughput Comparisons
Table I and Figure 6 compare the inference accuracy and token throughput of U-HLM with other inference methods. First, Table I shows that U-HLM outperforms both SLM and Rand-HLM in terms of cosine similarity across all datasets. In particular, with the Alpaca dataset, while Rand-HLM improves cosine similarity by only 2.08% compared to SLM, U-HLM achieves a 5.08% improvement, roughly 244% the accuracy gain. On average, U-HLM achieves up to 97.54% and 100.05% of the inference accuracy of LLM and HLM, respectively. Furthermore, as shown in Figure 6 (right), U-HLM yields the highest token throughput across all SNRs, aside from SLM. This advantage arises from U-HLM’s ability to accurately and selectively skip uplinks for tokens while maintaining performance, as evidenced by its high TSR and low TR in Figure 6 (left). In particular, the combination of a lower TR and a higher TSR compared to Rand-HLM highlights the effectiveness of U-HLM. Furthermore, U-HLM shows notable improvements in TSR, TR, and token throughput, especially when temperature perturbation is applied and under low average SNR conditions.

### v2

@@ Abstract
This paper studies a hybrid language model (HLM) architecture that integrates a small language model (SLM) operating on a mobile device with a large language model (LLM) hosted at the base station (BS) of a wireless network. The HLM token generation process follows the speculative inference principle: the SLM’s vocabulary distribution is uploaded to the LLM, which either accepts or rejects it, with rejected tokens being resampled by the LLM. While this approach ensures alignment between the vocabulary distributions of the SLM and LLM, it suffers from low token throughput due to uplink transmission and the computation costs of running both language models. To address this, we propose a novel HLM structure coined Uncertainty-aware opportunistic HLM (U-HLM), wherein the SLM locally measures its output uncertainty and skips both uplink transmissions and LLM operations for tokens that are likely to be accepted. This opportunistic skipping is enabled by our empirical finding of a linear correlation between the SLM’s uncertainty and the LLM’s rejection probability. We analytically derive the uncertainty threshold and evaluate its expected risk of rejection. Simulations show that U-HLM reduces uplink transmissions and LLM computations by 45.93%, while achieving up to 97.54% of the LLM’s inference accuracy and 2.54$\times$ faster token throughput than HLM without skipping.
@@ I Introduction
To this end, this paper seeks to find a solution for the HLM framework that improves token throughput while minimizing the degradation of inference accuracy. The approach builds on a straightforward yet effective strategy: skip uplink transmissions as well as LLM computations for tokens that are likely to be accepted. However, implementing this is challenging, as the SLM must independently predict whether a token will be accepted or rejected without access to the LLM’s vocabulary distribution. To enable this capability, we explore a method in which the SLM leverages uncertainty, a measure of the model’s self-assessed confidence in its outputs, to predict the LLM’s rejection probability. Among the various approaches we evaluate for measuring uncertainty in LLMs, our experiments reveal that uncertainty measured through temperature perturbation exhibits a strong linear correlation with the rejection probability in the HLM framework. Building on this observation, we propose a novel HLM framework called Uncertainty-aware opportunistic HLM (U-HLM). As illustrated in Figure 1, U-HLM enables the device to selectively transmit the SLM’s vocabulary distribution only for tokens with uncertainty exceeding a predefined threshold.
@@ III Uncertainty-Aware Hybrid Language Model for High-Throughput LLM Inference
In this section, we conduct a feasibility experiment inspired by the concept of uncertainty to investigate whether a device can predict the rejection probability of a token on its own without the server’s target probability. Based on this, we propose an Uncertainty-aware opportunistic Hybrid Language Model (U-HLM), in which the SLM measures the uncertainty and opportunistically skips uplink transmissions to the LLM if the estimated uncertainty does not exceed its threshold. Additionally, we design the uncertainty threshold for U-HLM based on the rejection probability threshold that determines when the device intends to skip transmissions. We also derive the expected rejection risk and its upper bound by incorporating the density of uncertainty into the analysis.
@@ IV-B Inference Accuracy and Token Throughput Comparisons
Table I and Figure 6 compare the inference accuracy and token throughput of U-HLM with other inference methods. Table I demonstrates that U-HLM outperforms both SLM and Rand-HLM in terms of cosine similarity across all datasets. For instance, with the Alpaca dataset, while Rand-HLM improves cosine similarity by only 2.08% over SLM, U-HLM achieves a 5.08% improvement—representing a 244% greater accuracy gain. On average, U-HLM achieves up to 97.54% and 100.05% of the inference accuracy of LLM and HLM, respectively. As shown in Figure 6 (right), U-HLM delivers the highest token throughput across all SNRs, second only to SLM, with a particularly notable 2.54$\times$ increase in token throughput compared to HLM without skipping. This advantage stems from U-HLM’s ability to accurately and frequently skip uplinks for tokens while maintaining inference quality. This is further evidenced by its higher TSR and lower TR compared to Rand-HLM in Figure 6 (left), reflecting a 45.93% reduction in uplink transmissions and LLM computations. Moreover, U-HLM demonstrates substantial improvements in TSR, TR, and token throughput, particularly when temperature perturbation is applied and under low average SNR conditions.
@@ V Concluding Remarks
Acknowledgments
This work was supported in part by the Korean Ministry of Science and ICT (MSIT) under the National Research Foundation of Korea (NRF) grant (No. 2023-11-1836), in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by MSIT (No. RS-2024-00404972), and in part by IITP under YKCS Open RAN Global Collaboration Center (IITP-2024-RS-2024-00434743) grant funded by MSIT.